Use the definition of limit to prove that the sequence { 1 √ n+1} converges to 0.


1
Expert's answer
2021-01-26T04:15:23-0500

"Let \space \epsilon >0"

We want "\mid\frac{1}{n^{0.5}}-0\mid<\epsilon\iff\frac{1}{n^{0.5}}<\epsilon\iff{n^{0.5}}>\frac{1}{\epsilon}"

So that "n>\frac{1}{{\epsilon}^{0.5}}"

Taking "N = \frac{1}{{\epsilon}^{0.5}}" we have n>N "\implies n>\frac{1}{{\epsilon}^{0.5}}" and for which we have

"\mid\frac{1}{n^{0.5}}-0\mid<\epsilon"

Meaning "\lim\limits_{x\to \infin}\frac{1}{n^{0.5}}=0" by the definition of limit